solve the system of linear equations:
4x + 10y = -28
6x + 7y = 6

Answer by unlockmath(1688)

Hello,
Let's multiply the first equation by 6 and the second one by -4. That will give us:
24x + 60y = -168
-24x -28y = -24
Now we can add both of these together and gives us:
32y=-192 Divide both sides by 32 will be:
y=-6 Now we can plug this into either original equation:
4x + 10(-6) = -28 Add 60 to both sides will be:
4x=32 Divide both sides by 4 to be:
x=8
Now plug both answers x=8 and y=-6 into both equations to see if it works out.
RJ
Check out a new book I wrote at:
www.math-unlock.com